OpenStudy (love_to_love_you):

Next question @Loser66

OpenStudy (love_to_love_you):

I'm gonna go through my steps, it just helps me think. Can you just lemme know if you see a problem?

OpenStudy (loser66):

Go ahead.

OpenStudy (love_to_love_you):

So for f(x) = -1/5x - 1 Switch roles: -1/5x = f^-1(x) - 1 Do I add 1 on both sides?

OpenStudy (love_to_love_you):

Sorry switching roles would be 1/5x = -f^-1(x) - 1

OpenStudy (loser66):

Nope, just switch the role for x and f(x) only, the numbers stay!!

OpenStudy (love_to_love_you):

ohhhh, so what would that be? sorry I suckkkk at math.

OpenStudy (loser66):

f(x) =(-1/5) x-1 step 1, switch the roles \(x=(-1/5) f^{-1}(x) -1 \)

OpenStudy (loser66):

ok? Numbers stay as they are.

OpenStudy (love_to_love_you):

Oh okay.

OpenStudy (love_to_love_you):

So now do I add 1 to each side?

OpenStudy (loser66):

yup

OpenStudy (love_to_love_you):

So 1 + x = (-1/5)f^-1(x) I suppose I do something with the -1/5 right?

OpenStudy (love_to_love_you):

multiply both sides by 5?

OpenStudy (loser66):

(-5) not 5

OpenStudy (love_to_love_you):

ok so -5 - 5x = f^-1(x)?

OpenStudy (loser66):

Bingo

OpenStudy (love_to_love_you):

so f^-1(x) = -5x -5 go on to graphing, do I just do the same thing you did with the whole -5(0) - 5 = -5 -5(1) - 5 = 0?

OpenStudy (loser66):

yes.

OpenStudy (love_to_love_you):

I feel like I did something wrong with the graphing, so would the two points be (0,-5) and (1,0)?

OpenStudy (loser66):

You are perfectly correct.

OpenStudy (love_to_love_you):



OpenStudy (loser66):

good

OpenStudy (love_to_love_you):

When I have mathway.com graph it for me it shows something different

OpenStudy (love_to_love_you):

The points on their graph are (-5, 0) and (-1, 0)

OpenStudy (loser66):

oh, yes!! my bad

OpenStudy (loser66):

(-5)*1 -5 = -10 not 0 but if you use (-5)*(-1) -5 =0

OpenStudy (loser66):

ha!! you see. I am wrong also. So, no need to feel bad on math.

OpenStudy (love_to_love_you):

so one of the points would be (1, -10)?

OpenStudy (loser66):

yes. :)

OpenStudy (love_to_love_you):

Okay, thank you so much! Could you possibly check my answer to the next question just to make sure I have it down? I won't need to go over the steps, just show you my answer when I get it

OpenStudy (loser66):

sure